A gas jar was filled with 150cm3 of hydrogen and100cm3 of Chlorine at room temperature,the two gases reacted in bright sunlight.After reaction and cooling the remaining gas mixture occupied the same volume . The mixture was bubbled through water , 100cm3 dissolved to produce a solution X and 50cm3 of a gas Y remained
(a) write equation for the reaction between hydrogen and chlorine
(b) what precautions must be taken when bubbling the product other this reaction into water
(c) identify gas Y
(d) what is the name given to solution X
(e)how will solution X react with
(I)CaCO3
(II)Mg ribbon
(III) AgNO3
1
Expert's answer
2020-06-03T13:34:02-0400

Solution: we +write the equation of interaction of hydrogen and chlorine under conditions when their mixture is illuminated: H2 + Cl2 = 2HCl. Moreover, we see that all the substances involved in the reaction are gaseous, then the reaction is calculated according to the volumes of the substances entering into the reaction. Then we have a gaseous mixture with a total volume of: 150cm3+100cm3=250cm3.At the same time, one volume of chlorine reacts to one volume of hydrogen. Then on 100cm3 chlorine have to 100cm3 hydrogen, and will remain unreacted: 150cm3-100cm3-50cm3 hydrogen. Then 2 * 100 cm3 = 200 cm3 of hydrogen chloride is formed. Then the volume, as written in the condition, remains unchanged: 50cm3 is volume of hydrogen and 200cm3 is volume of hydrogen chloride.

Further, passing the reaction gas through water, we get that part of it in the form of hydrogen chloride dissolves in water, and the hydrogen remaining after the reaction does not interact with it and does not dissolve in it. We see that its volume is 50 cm3, then it is gas Y.

Next, we pass the gas mixture through water. In this latter, hydrogen chloride dissolves, giving hydrochloric acid. At the same time, it follows from the condition that we dissolve 100 cm3 of hydrogen chloride in water, and we have formed 200 cm3 of hydrogen chloride. Then we get that in our case two portions of the solution are formed, in which 100 cm3 of hydrogen chloride are dissolved, and 50 cm3 of hydrogen or gas remains Y.

Thus, we got X - a solution of hydrogen chloride in water, or hydrochloric acid.

Hydrochloric acid will react with calcium carbonate by the reaction:

2HCl+CaCO3 =CaCl2+H2O.

The reaction of hydrochloric acid with magnesium tape is written as follows:

2HCl+Mg=MgCl2+H2.

Silver nitrate reaction:

HCl+AgNO3=AgCl+HNO3.

We see that in the first reaction a gas is released - carbon dioxide; in the second reaction, hydrogen gas is released; in the third reaction precipitates - silver chloride.


Answer: a) H2 + Cl2 = 2HCl

b)The process of dissolving hydrogen chloride in water is exothermic. Among other things, we pass a mixture of hydrogen chloride and hydrogen through water. Then, when this mixture is passed through, a large heating occurs, hydrogen begins, as a result, to leave the solution faster, and combines with the oxygen of the air, interacting with it and giving an explosion. In this regard, the main precaution should be taken to cool the vessel in which the dissolution of hydrogen chloride occurs.

с)Y is hydrogen, H2..

d)X is hydrochloric acid.

e) i)2HCl+CaCO3 =CaCl2+H2O;

ii)2HCl+Mg=MgCl2+H2;

iii)HCl+AgNO3=AgCl+HNO3.